At the present time nick is one-third as old as Karl. In 12 years, Karl's age will be six years less than twice nicks's age at the time. Find the present ages for both nick and karl
Found 2 solutions by stanbon, advanced_Learner:
Answer by stanbon(75887)

At the present time nick is one-third as old as Karl. In 12 years, Karl's age will be six years less than twice nicks's age at the time. Find the present ages for both nick and karl
-----
K = 3*N
K+12 = 3(N+12)-6
------------------
Substitute for "K" and solve for "N"::
3N+12 = 3N + 30
----
That is a contradiction.
----
Ans:: No solution.
Cheers,
Stan H.

Answer by advanced_Learner(501)

At the present time nick is one-third as old as Karl. In 12 years, Karl's age will be six years less than twice nicks's age at the time. Find the present ages for both nick and karl

K=3*N
K+12=2(N+12)-6
N=(6)
SO
K=18
just
in 12 years
N=18
K=30 which is 6 years less than 36
checks!